OpenStudy (anonymous):

In a direct variation, y is negative six when x is twenty-four. What is y when x equals sixteen? negative sixty-four negative fourteen negative four negative two

OpenStudy (anonymous):

Negative four because 24/6=4/1 4x4=16 therefore the denominator would be four or 16/4~~4/1 :)